OpenStudy (anonymous):

2x + y = 7 5x + y = 9 Solve the system of equations. (-2/3, 25/3) (2/3, 17/3) (2/3, 25/3)

OpenStudy (luffingsails):

Try multiplying the first equation by -1... then adding the result to the second equation.

OpenStudy (anonymous):

What?

OpenStudy (luffingsails):

You have a system of equations. One of the methods to solve them is by manipulation. So, multiply eq 1 by -1.. like this: -1(2x+y) = -1(7) -2x - y = -1 Now add that to the second equation -2x - y = -1 5x + y = 9 3x = 8 x = 8/3 Now substitute the x value into either equation to get y.

OpenStudy (anonymous):

8/3 isn't one of the answer choices.

OpenStudy (luffingsails):

Sorry, I typed too quick.. -1(2x + y) = -1(7) -2x - y = -7 Adding -2x - y = -7 5x + y = 9 results in: 3x = 2 x = 2/3 Then substitute. Very sorry!!!

OpenStudy (anonymous):

No worries:)